# Geological Everything Model 3D: A Promptable Foundation Model for Unified and Zero-Shot Subsurface Understanding

### v1

@@ 2.3.3 Emergent structural interpretation of unseen Martian SHARAD radagrams
We evaluated GEM’s capability to interpret complex internal reflectors within the Martian polar ice caps using 3D SHARAD radar data. These 3D datasets, derived from thousands of orbital radar soundings, reveal the intricate stratigraphy of Planum Boreum and Planum Australe, which record layered deposition, unconformities, and possible impact-related structures accumulated over millions of years of climate cycling on Mars. However, interpreting these internal layers is highly challenging due to limited vertical resolution, low signal-to-noise ratio, and frequent clutter arising from surface topography and ionospheric distortions. Critically, there are currently no effective automated methods for extracting internal radar reflectors, and manual interpretation remains the norm—an inefficient and subjective process that limits scientific inference..
@@ 2.3.4 Fully stratigraphic interpretation
In addition to structural interpretation, GEM also generalizes well to interpret unconformities (cyan surfaces in Figure 3 (e, f)) and estimate a RGT volume for achieving fully stratigraphic interpretation[50, 19]. By constructing a stratigraphic time field that represents isochronous surfaces throughout the seismic volume, RGT captures detailed stratification, unconformities, and complex structural features, while effectively encoding information about depositional sequences and geological structures. Compared with traditional discrete horizon picking methods, RGT offers a dense, coherent, and geologically consistent representation, serving as a unified foundation for tasks such as stratigraphic correlation, sedimentary analysis.
@@ 2.4 Geobody Segmentation
The synthetic test set for channels was derived from the CIGChannel [51] dataset published by Wang et al. and consists of 100 samples. The synthetic test set for karsts was taken from the KarstSeg [52] dataset released by Wu et al., comprising 20 samples. For real data validation, 30 two-dimensional slices were randomly selected from the Parihaka and Romney (Figure 10 (b-1)) datasets for channel evaluation. The baseline methods used for comparison experiments include UNet, DeepLabV3, and HRNet.
@@ 2.4.2 Karst segmentation
Figure 4 (d) illustrates the zero-shot application of GEM for karst geobody interpretation in the Fort Worth Basin, a hydrocarbon-rich foreland basin in North Central Texas. This region is geologically characterized by extensive paleokarst development within the Ellenburger carbonate formation, formed during prolonged subaerial exposure from the Late Ordovician to Early Pennsylvanian. Subsequent burial and compaction induced collapse structures, resulting in complex karst systems that appear as distinct geomorphological features in seismic data.
@@ 2.5 Well-Log Prompted Geophysical Property Modeling
To demonstrate GEM’s modeling capabilities, we compare its performance against three conventional AI models trained under supervised, multidimensional inversion frameworks. While these baselines rely on paired seismic–well data and task-specific training, GEM performs inference directly using well logs as prompts, without any retraining. In the quantitative results, for the synthetic dataset, where a complete 3D velocity volume is available, we report both semantic-level and patch-level performance metrics, including Fr´echet Inception Distance (FID) [55] and Structural Similarity Index (SSIM) [56]. For the field datasets, that lack full-volume ground truth, evaluation is conducted using pixel-level Mean Absolute Error (MAE) calculated at blind well locations—that is, wells excluded from both the training set of conventional models and the prompt input to GEM. This ensures a fair and consistent comparison of generalization performance across methods. The quantitative metrics in Figure 5a clearly demonstrate GEM’s superior performance in property modeling: it achieves consistently higher SSIM and lower MAE on synthetic data, and lower MAE on field datasets, outperforming conventional supervised models across all test cases
@@ 2.5.2 Acoustic impedance modeling on the Teapot Dome
In Figure 5 (c), GEM demonstrates superior performance in capturing thinlayer structures with greater clarity and continuity. It effectively suppresses the vertical ringing and lateral instability commonly observed in traditional methods. While maintaining consistency with the seismic structural framework, the impedance volumes generated by GEM exhibit higher spatial resolution and stronger stratigraphic continuity, particularly in zones with thin interbedded sandstone and mudstone. Figure 5 (c-3) presents the blind-well test results, demonstrating a high degree of alignment between GEM’s predictions and the impedance trends observed in well logs. The corresponding MAE values further confirm the model’s accuracy. In addition, Appendices Figure 11 (a) provides a qualitative comparison of property modeling results between GEM and several baseline methods. GEM consistently outperforms the others, particularly in terms of vertical resolution and the preservation of geological structures.
@@ 2.5.3 Multi-property modeling on the Netherlands F3
The Netherlands F3 dataset is a widely used benchmark for 3D seismic interpretation and property modeling. Geologically, it features multi-phase fault systems, progradational slope deposits, and a regionally extensive angular unconformity that separates genetically distinct stratigraphic packages. These structures present strong phase shifts, impedance discontinuities, and highamplitude reflectors—making seismic inversion in this dataset particularly challenging. Such complex features demand models capable of capturing nonlinear responses and maintaining stratigraphic coherence across discontinuities.
@@ 3.2 Self Supervised Pre-training
We adopt a BERT-style pretraining strategy for convolutional networks, specifically leveraging the Sparse Masked Modeling (SparK) approach [58], with tailored modifications for the GEM backbone. During the masked image modeling (MIM) process in 3D data, we employ a higher masking ratio of 80%, exceeding the 60% used in the original SparK method. This adjustment is motivated by the inherent characteristics of seismic data, where neighboring slices exhibit high similarity and repeated textural patterns. As a result, reconstructing 3D seismic volumes is considerably less challenging than reconstructing natural 2D images. Increasing the reconstruction difficulty encourages the model to develop more generalizable representations, thereby improving performance in downstream tasks.
@@ 3.3 Structure-aware Perceptual Network
Training is conducted on NVIDIA H20 GPUs with a batch size of , using the AdamW optimizer with a learning rate of , , , and a total of 300,000 iterations.
@@ 3.4 Supervised Fine-tuning with Mixed Prompting
Due to the ambiguity and anisotropy inherent in subsurface imaging data, which often result in non-uniqueness, practical applications tend to prioritize the reliability of a solution among all possible outcomes rather than its absolute correctness. This consideration is particularly critical in tasks such as structural modeling, RGT estimation, and inversion, where the objective is to recover geologically plausible structures from incomplete information. Under such circumstances, only relative correctness can be meaningfully defined. To address this challenge, we adopt a relativistic GAN strategy [60]. This approach shifts the modeling objective away from generating a single correct solution and instead encourages the model to learn how to select relatively better and geologically more credible structures from among multiple plausible solutions. This modeling paradigm aligns more naturally with the inherent uncertainty and non-uniqueness of seismic data. The corresponding formulation is given as follows,
@@ 4 Conclusion
We introduce GEM, a prompt-based foundation model for subsurface imaging that unifies structural interpretation, stratigraphic analysis, geobody segmentation, and property modeling within a single generative framework. By reformulating geophysical tasks as prompt-guided conditional generation over structurally inferred latent spaces, GEM supports flexible human interaction and generalizes across surveys, data types, and geological settings without task-specific retraining.
@@
To our knowledge, GEM is the first generative foundation model in geosciences that supports expert-in-the-loop interaction, prompt-guided reasoning, and zero-shot generalization across tasks, surveys, and data types. More broadly, this work establishes a new modeling paradigm for scientific AI in domains where data is structurally rich but semantically sparse, and where insight emerges from contextual, expert-informed interpretation. GEM exemplifies how generative models can be steered by structural priors, enabling interpretable, scalable, and collaborative modeling of the Earth’s interior. We envision this structure-aware, promptable approach extending to broader domains in scientific machine learning, including physical simulation, inverse problems, and multi-modal Earth system modeling.
@@ Appendix E Backbone Network Architecture
Appendix F GUI for Interaction
Figure 13 displays the graphical user interface (GUI) developed for interactive structural interpretation and geobody segmentation. The repository provides examples demonstrating how this interface is used for interactive interpretation.

### v2

@@ 2.3.3 Emergent structural interpretation of unseen Martian SHARAD radagrams
We evaluated GEM’s capability to interpret complex internal reflectors within the Martian polar ice caps using 3D SHARAD radar data. These 3D datasets, derived from thousands of orbital radar soundings, reveal the intricate stratigraphy of Planum Boreum and Planum Australe, which record layered deposition, unconformities, and possible impact-related structures accumulated over millions of years of climate cycling on Mars. However, interpreting these internal layers is highly challenging due to limited vertical resolution, low signal-to-noise ratio, and frequent clutter arising from surface topography and ionospheric distortions. Critically, there are currently no effective automated methods for extracting internal radar reflectors, and manual interpretation remains the norm—an inefficient and subjective process that limits scientific inference.
@@ 2.3.4 Fully stratigraphic interpretation
In addition to structural interpretation, GEM also generalizes well to interpret unconformities (cyan surfaces in Figure 3 (e, f)) and estimate a RGT volume for achieving fully stratigraphic interpretation[53, 19]. By constructing a stratigraphic time field that represents isochronous surfaces throughout the seismic volume, RGT captures detailed stratification, unconformities, and complex structural features, while effectively encoding information about depositional sequences and geological structures. Compared with traditional discrete horizon picking methods, RGT offers a dense, coherent, and geologically consistent representation, serving as a unified foundation for tasks such as stratigraphic interpretation, sedimentary analysis.
@@ 2.4 Geobody Segmentation
The synthetic test set for channels was derived from the CIGChannel [54] dataset with consists of 100 samples. The synthetic test set for karsts was taken from the KarstSeg [55] dataset, comprising 20 samples. For real data validation, 30 two-dimensional slices were randomly selected from the Parihaka and Romney (Figure 10 (b-1)) datasets and manually labeled for channel evaluation. The baseline methods used for comparison experiments include UNet, DeepLabV3, and HRNet.
@@ 2.4.2 Karst segmentation
Figure 4 (d) illustrates the application of GEM for karst geobody interpretation in the Fort Worth Basin, a hydrocarbon-rich foreland basin in North Central Texas. This region is geologically characterized by extensive paleokarst development within the Ellenburger carbonate formation, formed during prolonged subaerial exposure from the Late Ordovician to Early Pennsylvanian. Subsequent burial and compaction induced collapse structures, resulting in complex karst systems that appear as distinct geomorphological features in seismic data.
@@ 2.5 Well-Log Prompted Geophysical Property Modeling
To demonstrate GEM’s modeling capabilities, we compare its performance against three conventional AI models trained under supervised, multidimensional inversion frameworks. While these baselines rely on paired seismic–well data and task-specific training, GEM performs inference directly using well logs as prompts, without any retraining. In the quantitative results, for the synthetic dataset, where a complete 3D velocity volume is available, we report both semantic-level and patch-level performance metrics, including Frechet Inception Distance (FID) [59] and Structural Similarity Index (SSIM) [60]. For the field datasets, that lack full-volume ground truth, evaluation is conducted using pixel-level Mean Absolute Error (MAE) calculated at blind well locations—that is, wells excluded from both the training set of conventional models and the prompt input to GEM. This ensures a fair and consistent comparison of generalization performance across methods. The quantitative metrics in Figure 5a clearly demonstrate GEM’s superior performance in property modeling: it achieves consistently higher SSIM and lower MAE on synthetic data, and lower MAE on field datasets, outperforming conventional supervised models across all test cases.
@@ 2.5.2 Acoustic impedance modeling on the Teapot Dome
In Figure 5 (c), GEM demonstrates superior performance in capturing thinlayer structures with greater clarity and continuity. It effectively suppresses the vertical ringing and lateral instability commonly observed in traditional methods (Figure 11). While maintaining consistency with the seismic structural framework, the impedance volumes generated by GEM exhibit higher spatial resolution and stronger stratigraphic continuity, particularly in zones with thin interbedded sandstone and mudstone. Figure 5 (c-3) presents the blind-well test results, demonstrating a high degree of alignment between GEM’s predictions and the impedance trends observed in well logs. The corresponding MAE values further confirm the model’s accuracy. In addition, Appendices Figure 11 (a) provides a qualitative comparison of property modeling results between GEM and several baseline methods. GEM consistently outperforms the others, particularly in terms of vertical resolution and the preservation of geological structures.
@@ 2.5.3 Multi-property modeling on the Netherlands F3
The Netherlands F3 dataset is a widely used benchmark for 3D seismic interpretation and property modeling. Geologically, it features multi-phase fault systems, progradational slope deposits, and a regionally extensive angular unconformity that separates genetically distinct stratigraphic packages. These structures present strong phase shifts, impedance discontinuities, and highamplitude reflectors—making property modeling in this dataset particularly challenging. Such complex features demand models capable of capturing nonlinear responses and maintaining stratigraphic coherence across discontinuities.
@@ 3.2 Self Supervised Pre-training
We adopt a BERT-style pretraining strategy for convolutional networks, specifically leveraging the Sparse Masked Modeling (SparK) approach [62], with tailored modifications for the GEM backbone. During the masked image modeling (MIM) process in 3D data, we employ a higher masking ratio of 80%, exceeding the 60% used in the original SparK method. This adjustment is motivated by the inherent characteristics of seismic data, where neighboring slices exhibit high similarity and repeated textural patterns. Therefore, under the same missing ratio, reconstructing 3D seismic data is easier than reconstructing 2D images. Increasing the reconstruction difficulty encourages the model to develop more generalizable representations, thereby improving performance in downstream tasks.
@@ 3.3 Structure-aware Perceptual Network
Training of the perceptual network is conducted on NVIDIA H20 GPUs with a batch size of , using the AdamW optimizer with a learning rate of , , , and a total of 300,000 iterations.
@@ 3.4 Supervised Fine-tuning with Mixed Prompting
Due to the ambiguity and anisotropy inherent in subsurface imaging data, which often result in non-uniqueness, practical applications tend to prioritize the reliability of a solution among all possible outcomes rather than its absolute correctness. This consideration is particularly critical in tasks such as structural modeling, RGT estimation, and property modeling, where the objective is to recover geologically plausible structures from incomplete information. Under such circumstances, only relative correctness can be meaningfully defined. To address this challenge, we adopt a relativistic GAN strategy [64]. This approach shifts the modeling objective away from generating a single correct solution and instead encourages the model to learn how to select relatively better and geologically more credible structures from among multiple plausible solutions. This modeling paradigm aligns more naturally with the inherent uncertainty and non-uniqueness of seismic data. The corresponding formulation is given as follows,
@@ 4 Conclusion
We introduce GEM, a prompt-based foundation model for subsurface imaging that unifies structural interpretation, stratigraphic analysis, geobody segmentation, and property modeling within a single generative framework. By reformulating all these tasks as prompt-guided conditional generation over structurally inferred latent spaces, GEM supports flexible human interaction and generalizes across surveys, data types, and geological settings without task-specific retraining.
@@
To our knowledge, GEM is the first generative foundation model in subsurface understanding that supports expert-in-the-loop interaction, prompt-guided reasoning, and zero-shot generalization across tasks, surveys, and data types. More broadly, this work establishes a new modeling paradigm for scientific AI in domains where data is structurally rich but semantically sparse, and where insight emerges from contextual, expert-informed interpretation. GEM exemplifies how generative models can be steered by structural priors, enabling interpretable, scalable, and collaborative modeling of the Earth’s interior. We envision this structure-aware, promptable approach extending to broader domains in scientific machine learning, including physical simulation, inverse problems, and multi-modal Earth system modeling.
@@ Appendix E Backbone Network Architecture
Appendix F Training and Inference Cost
All training stages were carried out using eight NVIDIA H20 GPUs (96 GB each). The model comprises 0.67 billion parameters. During the self-supervised pretraining phase, the input size was with a batch size of 64. The model was trained for 500,000 steps, each taking approximately 1.8 seconds, totaling around 10 days. For downstream tasks, the first 300,000 steps used an input size of with a batch size of 8, and each step took approximately 1.7 seconds. The remaining 200,000 steps were trained under PyTorch’s checkpoint mode with an input size of and a batch size of 8, requiring about 3.2 seconds per step, for a total of roughly 13 days. During inference, processing a volume requires approximately GFLOPs. At float16 precision, the memory usage is around 30 GB.
For interactive inference tasks such as structural interpretation and geobody segmentation, we recommend resizing the data to lower resolutions or smaller volumes on the GPU to improve interactivity. In contrast, tasks like property modeling and RGT estimation benefit from inference at the original resolution or volume size, which can be efficiently handled on CPU memory (CPU memory is more accessible and easier to scale, allowing for inference on larger volumes). This sacrifices inference speed but yields higher voxel-level accuracy and resolution.
Appendix G GUI for Interaction
Figure 13 shows the interactive GUI, more examples are provided on the accompanying project page.